OpenStudy (anonymous):

Which property would be used to isolate the variable h in this equation? h + 12 = 26

OpenStudy (anonymous):

A. addition property of equality B. commutative property of addition C. multiplication property of equality D. division property of equality

OpenStudy (anonymous):

Is commutative the same as subtraction?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

smh hold on then

OpenStudy (anonymous):

okay girl

OpenStudy (anonymous):

That's the only choices huh? Because I still think it would be the subtraction one unless you multiply both sides to get rid of the 12?

OpenStudy (anonymous):

those are the only choices omg

OpenStudy (anonymous):

I say C since you can multiply (1/12)(h+12) = (1/12)(26) buuuuut that's just a guess

OpenStudy (anonymous):

thanks

Directrix (directrix):

@nofkfmfmfmfmfmfmfmf @RavenLynette Think about addition (option A) as an answer. Adding negative numbers is the equivalent of subtraction. Do you agree? h + 12 = 26 h + 12 + (-12) = 26 + (-12) h = ?

OpenStudy (anonymous):

That does make sense too and it's probably the better answer since it doesn't give you a fraction, I hadn't thought of it that way, thanks

OpenStudy (anonymous):

multiplication property